### v1

@@ 5.2 Evaluation in Simulated Environment
Figure 5: Comparison between minimum snap and adapted trajectories in response to waypoint deviations, where waypoints are shifted by 2 meters and rotated by 30 degrees. In the first row, initial and updated waypoints are shown alongside baseline and trained policy trajectories. The second figure in the first row illustrates the reference motor speed commands equally sampled from each trajectory segment. The third figure demonstrates simulated motor speeds, revealing that the adapted trajectory is 9.22% faster than the baseline while maintaining motor speeds within the admissible range. The second row compares tracking errors, where the adapted trajectory remains below the threshold, unlike the minimum snap trajectory. The last row presents speed profiles before and after waypoint deviations. Background feasibility colormap is generated by random speed profile sampling and averaging of reference motor speed feasibility. Both trajectories are initially within the feasible region (Initial speed profile), but as waypoints change, the MFRL policy keeps the trajectory feasible, while the MinSnap baseline ventures into infeasible regions (Adapted speed profile).
@@
To evaluate our policy, we reconstruct trajectories starting from the midpoint of the waypoint sequences. Subsequently, we introduce random shifts and rotations to the remaining waypoints to assess whether our policy could generate feasible trajectories. The testing dataset is derived from 4,000 randomly generated waypoint sequences, with trajectory times ($T^{\text{MS}}$) determined through numerical simulations. The testing trajectories consist of an average speed of 3.9 $\mathrm{m}\mathrm{/}\mathrm{s}$, with a maximum of 11.2 $\mathrm{m}\mathrm{/}\mathrm{s}$, and an average acceleration of 0.7 g, with a maximum of 2.7 g, where g denotes standard gravity acceleration ($\text{g}=9.81$\mathrm{m}\mathrm{/}\mathrm{s}^{2}$$).
@@
Table 2 presents a comparison of tracking error between the baseline minimum snap (MinSnap) trajectory and trajectories generated by our trained policy (MFRL). When there are deviations in the position and yaw of the remaining waypoints, the baseline technique employs the original trajectory’s time allocation and solves the minimum-snap quadratic programming problem using the updated waypoint positions to adapt the trajectory. As the remaining waypoints deviate further, the baseline method encounters greater average tracking error as well as increased error variance, reflecting its struggle to track a larger portion of the trajectory. On the other hand, the trained policy adjusts the time allocations to generate a trajectory that consistently maintains acceptable tracking errors. The table also showcases the time reduction ($\Delta$ Time) achieved by our trained policy, where a positive reduction indicates faster trajectories compared to the baseline method. Initial time allocations are optimized for smooth trajectories; thus, adapting these allocations demands extra time in response to random waypoint deviations. Consequently, time reduction is less in such scenarios compared to non-deviation cases. Even so, the trained model consistently achieves faster trajectories than the baseline. Figure 5 compares motor speed and tracking error between the baseline method and the trained policy on an example trajectory with randomly deviated waypoints. As a response to these deviations, the trained policy updates the trajectory to maintain an admissible tracking error and reduce flight time, unlike the baseline method, which fails to generate a feasible trajectory.
@@
[…]
Table 4: Comparison of tracking error and trajectory time between the minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL). The policy is trained with the first and second fidelity levels in the large target space $L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$]$. The trained policy generates trajectories that are 6.23% faster than the baseline when waypoints are unchanged.
@@
[…]
In Table 3, we compare our approach with two widely used methods: 1) MinSnap Heuristic: a heuristic method that adjusts time allocation proportional to the square root of the trajectory length and 2) Supervised-only: a policy model trained solely with supervised learning. The pretrained model, utilized for MFRL initialization, is used to evaluate the second method, employing only the time allocation output with the identity smoothness weight. The first method fails to adapt the trajectory, leading to diverging tracking errors when waypoints are randomly shifted. In contrast, the second method can update the trajectory to remain feasible, yet it results in a longer trajectory time compared to the reinforcement learning-trained model.
We further apply the training in a larger space ($L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$]$) using scaled waypoint sequence data. In the wider space, trajectories could attain higher linear speeds, resulting in testing trajectories with an average speed of 5.5 $\mathrm{m}\mathrm{/}\mathrm{s}$ and a maximum of 15.1 $\mathrm{m}\mathrm{/}\mathrm{s}$, as well as an average acceleration of 0.7 g and a maximum of 2.9 g. As shown in Table 4, our trained model effectively updates trajectories while maintaining tracking accuracy, achieving even greater time reduction compared to the smaller space. In smaller spaces, tracking error bounds are typically determined by yaw tracking error due to the need for agile turns in confined areas. In contrast, in larger spaces, position tracking error becomes more prominent, as there is ample room for linear acceleration and reaching the vehicle’s maximum speed, providing our MFRL model with additional opportunities for optimization.
@@ 5.3 Analysis of MFRL Policy
Comparing the utilization factor between baseline trajectories and those from the trained model in Figure 7 reveals that the trained policy reduces the variance of the utilization factor and increases its mean. This indicates that MFRL trains the policy to maintain motor utilization at an optimal level, efficiently utilizing the vehicle’s capacity while ensuring feasibility. As shown in the right plot of Figure 7, when the initial minimum snap trajectory has low motor utilization, the trained model achieves significantly greater reductions (about 12 % faster) in flight time. Furthermore, the time reduction’s variation associated with motor utilization is smaller compared to other trajectory features, highlighting their strong correlation. Table 5, presenting data from the default space size, and Table 6, from a larger space, compare parameters such as speed, acceleration, and motor utilization between minimum-snap and MFRL trajectories. Both tables reveal that, while average speeds and accelerations are comparable, MFRL trajectories show higher maximums and a significantly greater average motor utilization with lower variance.
Table 5: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL) ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$]$).
@@
Table 6: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL) ($L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$]$).
@@ 5.4 Real-World Experiment
[…]
@@
We further evaluate the proposed algorithm by training the planning policy in a real-world scenario and testing the policy on a dataset of 50 randomly generated waypoint sequences, each with trajectory times $T^{\text{MS}}$ determined from real-world flight experiments. The reward estimator is trained using three different fidelity levels, including real-world flight experiments. Table 7 presents a comparison of the tracking error and the time reduction ($\Delta$ Time) achieved by policy models trained in simulated environments (using two fidelity levels) and real-world scenarios (utilizing three fidelity levels). Notably, when waypoints deviate, both the models trained with MFRL in simulation and real-world experiments adapt the trajectory better than the baseline method. The model trained with real-world experiments further optimizes the trajectory to make it even faster than the policy only trained in simulation. Table 9 compares speed, acceleration, and motor utilization between the baseline trajectories and the trajectories from the trained policies. Similar to the simulated results, while average speeds and accelerations are comparable, MFRL trajectories show higher maximums and a significantly greater average motor utilization.
Table 9: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) trajectories generated by the policy trained in simulation (MFRL (sim)), and trajectories generated by the policy trained with real-world experiments (MFRL (real)) ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$]$).
$v_{\text{avg}}$ $v_{\text{max}}$ $a_{\text{avg}}$ $a_{\text{max}}$ $U_{\text{motor}}$
@@
Additionally, we assess the trained model’s performance in an online planning scenario, where dynamically shifting waypoints necessitate real-time trajectory adaptation within the environment shown in Figure 11. We put the waypoints in the middle of two gates and continuously moving the gates to update the position of the waypoints. We commit a portion of the first trajectory segment, taking into account inference and communication delays, while adjusting the time allocation for the remainder. The time allocation between waypoints is determined by the trained policy, and we obtain the actual polynomial trajectory through quadratic programming. The trained policy is executed on the Titan Xp GPU in the host computer and communicates with the vehicle’s microcontroller to update the trajectory. By compiling the model with TensorRT, we significantly enhance the inference speed. This enables our trained model to adjust the trajectory in under 2 milliseconds, allowing real-time trajectory adaptation—a marked contrast to the baseline minimum-snap method, which takes several minutes for trajectory generation. Since the planning policy comprises simple MLPs, the inference time remains consistent even on embedded GPUs, averaging 3 milliseconds when executed on Jetson Orin. Table 8 presents a comparison of inference times between the baseline method and our trained model. The baseline method incurs significant computation time due to its line search procedure, particularly during real-world flight experiments where evaluating a trajectory with multiple speed profiles can take several minutes. Our proposed method streamlines this by substituting the line search with policy inference, making it suitable for real-time applications. Figure 12 compares the inference times based on the number of waypoints. Even with a small number of waypoints, the policy inference time is substantially lower, requiring far less computation time. The supplementary video includes demonstrations of the real-world experimental results.

### v2

@@ 5.2 Evaluation in Simulated Environment
Figure 11: Comparison between minimum snap and adapted trajectories in response to waypoint deviations, where waypoints are shifted by 2 meters and rotated by 30 degrees. (a) The initial and updated waypoints are shown alongside baseline and trained policy trajectories. (b) Reference motor speed commands equally sampled from each trajectory segment. (c) Simulated motor speeds, demonstrating that the adapted trajectory is 9.22% faster than the baseline while maintaining motor speeds within the admissible range. (d) and (e) Comparison of tracking errors, where the adapted trajectory remains below the threshold, unlike the minimum snap trajectory. The last row, (f) and (g) presents speed profiles before and after waypoint deviations. Background feasibility colormap is generated by random speed profile sampling and averaging of reference motor speed feasibility. Both trajectories are initially within the feasible region (f), but as waypoints change, the MFRL policy keeps the trajectory feasible, while the MinSnap baseline ventures into infeasible regions (g).
@@
To evaluate our policy, we reconstruct trajectories starting from the midpoint of the waypoint sequences. Subsequently, we introduce random shifts and rotations to the remaining waypoints to assess whether our policy could generate feasible trajectories. The testing dataset is derived from 4,000 randomly generated waypoint sequences, with trajectory times ($T^{\text{MS}}$) determined through numerical simulations. The testing trajectories consist of an average speed of 3.9 $\mathrm{m}\mathrm{/}\mathrm{s}$, with a maximum of 11.2 $\mathrm{m}\mathrm{/}\mathrm{s}$, and an average acceleration of 0.7 g, with a maximum of 2.7 g, where g denotes standard gravity acceleration ($\text{g}=9.81$\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$).
Table 3 presents a comparison between the baseline minimum snap (MinSnap) trajectory and trajectories generated by our trained policy (MFRL). Along with the time reduction, we compare the mean and standard deviation of tracking errors for trajectories. The position and yaw tracking errors, $\bar{r}^{\text{err}}$ and $\bar{\psi}^{\text{err}}$, are calculated as the maximum deviation between the vehicle’s pose and the reference pose:
$\bar{r}^{\text{err}}$ $=\max_{t}\>\lVert p_{r}(t)-r(t)\rVert,$
$\bar{\psi}^{\text{err}}$ $=\max_{t}\>|p_{\psi}(t)-\psi(t)|$ (33)
@@
where $r(t)$ and $\psi(t)$ are the reference position and yaw angle respectively, as defined in (16). $p_{r}(t)$ and $p_{\psi}(t)$ represent the simulated position and yaw angle of the vehicle at time $t$. For clarity, while constraining the tracking error maintains a marginal time difference between actual and reference trajectory times, this difference is negligible compared to our method’s overall time reduction. When there are deviations in the position and yaw of the remaining waypoints, the baseline technique employs the original trajectory’s time allocation and solves the minimum-snap quadratic programming problem using the updated waypoint positions to adapt the trajectory. As the remaining waypoints deviate further, the baseline method encounters greater average tracking error as well as increased error variance, reflecting its struggle to track a larger portion of the trajectory. On the other hand, the trained policy adjusts the time allocations to generate a trajectory that consistently maintains acceptable tracking errors. The table also showcases the time reduction ($\Delta$ Time) achieved by our trained policy, where a positive reduction indicates faster trajectories compared to the baseline method. Initial time allocations are optimized for smooth trajectories; thus, adapting these allocations demands extra time in response to random waypoint deviations. Consequently, time reduction is less in such scenarios compared to non-deviation cases. Even so, the trained model consistently achieves faster trajectories than the baseline. Figure 11 compares motor speed and tracking error between the baseline method and the trained policy on an example trajectory with randomly deviated waypoints. As a response to these deviations, the trained policy updates the trajectory to maintain an admissible tracking error and reduce flight time, unlike the baseline method, which fails to generate a feasible trajectory.
@@
0.0 $\mathrm{m}$ /   0 $\mathrm{\SIUnitSymbolDegree}$ 0.08 $\pm$ 0.02 $\mathrm{m}$ 12.9 $\pm$ 1.3 $\mathrm{\SIUnitSymbolDegree}$ 0.00 % 0.08 $\pm$ 0.01 $\mathrm{m}$ 13.2 $\pm$ 4.7 $\mathrm{\SIUnitSymbolDegree}$ -0.69 %
1.0 $\mathrm{m}$ / 15 $\mathrm{\SIUnitSymbolDegree}$ 0.17 $\pm$ 2.36 $\mathrm{m}$ 21.0 $\pm$ 27.9 $\mathrm{\SIUnitSymbolDegree}$ -0.50 % 0.09 $\pm$ 0.04 $\mathrm{m}$ 16.0 $\pm$ 16.6 $\mathrm{\SIUnitSymbolDegree}$ -1.15 %
2.0 $\mathrm{m}$ / 30 $\mathrm{\SIUnitSymbolDegree}$ 0.24 $\pm$ 2.39 $\mathrm{m}$ 39.7 $\pm$ 52.2 $\mathrm{\SIUnitSymbolDegree}$ -1.34 % 0.11 $\pm$ 0.09 $\mathrm{m}$ 24.1 $\pm$ 34.7 $\mathrm{\SIUnitSymbolDegree}$ -2.01 %
3.0 $\mathrm{m}$ / 45 $\mathrm{\SIUnitSymbolDegree}$ 0.39 $\pm$ 3.50 $\mathrm{m}$ 58.3 $\pm$ 65.3 $\mathrm{\SIUnitSymbolDegree}$ -2.14 % 0.13 $\pm$ 0.18 $\mathrm{m}$ 32.5 $\pm$ 46.2 $\mathrm{\SIUnitSymbolDegree}$ -3.12 %
Table 5: Comparison of tracking error and trajectory time between the minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL). The policy is trained with the first and second fidelity levels in the large target space $L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$\$]. The trained policy generates trajectories that are 6.23% faster than the baseline when waypoints are unchanged.
@@
0.0 $\mathrm{m}$ /   0 $\mathrm{\SIUnitSymbolDegree}$ 0.09 $\pm$ 0.02 $\mathrm{m}$ 13.4 $\pm$ 0.8 $\mathrm{\SIUnitSymbolDegree}$ 0.12 $\pm$ 0.56 $\mathrm{m}$ 12.9 $\pm$ 7.0 $\mathrm{\SIUnitSymbolDegree}$ 6.23 %
2.0 $\mathrm{m}$ / 15 $\mathrm{\SIUnitSymbolDegree}$ 0.14 $\pm$ 1.30 $\mathrm{m}$ 18.5 $\pm$ 21.4 $\mathrm{\SIUnitSymbolDegree}$ 0.13 $\pm$ 0.63 $\mathrm{m}$ 13.7 $\pm$ 8.8 $\mathrm{\SIUnitSymbolDegree}$ 5.74 %
4.0 $\mathrm{m}$ / 30 $\mathrm{\SIUnitSymbolDegree}$ 0.24 $\pm$ 2.16 $\mathrm{m}$ 29.3 $\pm$ 37.5 $\mathrm{\SIUnitSymbolDegree}$ 0.14 $\pm$ 0.87 $\mathrm{m}$ 16.8 $\pm$ 17.8 $\mathrm{\SIUnitSymbolDegree}$ 4.71 %
6.0 $\mathrm{m}$ / 45 $\mathrm{\SIUnitSymbolDegree}$ 0.46 $\pm$ 3.24 $\mathrm{m}$ 42.0 $\pm$ 50.0 $\mathrm{\SIUnitSymbolDegree}$ 0.15 $\pm$ 0.54 $\mathrm{m}$ 20.3 $\pm$ 25.0 $\mathrm{\SIUnitSymbolDegree}$ 4.49 %
The incorporation of the first fidelity level evaluations proves essential to our methodology. Their absence would necessitate a prohibitively large amount of the second fidelity evaluations to replace the current data volume, making collection impractically time-consuming. Alternatively, maintaining only the current limited quantity of the second fidelity evaluations without the first fidelity data would result in highly inaccurate feasibility constraint predictions, causing the MFRL training process to diverge.
In Table 4, we compare our approach with two widely used methods: 1) MinSnap Heuristic: a heuristic method that adjusts time allocation proportional to the trajectory length, chosen for its comparable computation speed:
$x_{t,i}=\frac{\lVert\tilde{p}_{\text{new}}^{i}-\tilde{p}_{\text{new}}^{i-1}\rVert}{\lVert\tilde{p}^{i}-\tilde{p}^{i-1}\rVert}\>x^{\text{MS}}_{t,i}$ (34)
and 2) Supervised-only: a policy model trained to imitate minimum snap optimization output, serving as a proxy for comparison with optimization-based methods. The pretrained model, utilized for MFRL initialization, is used to evaluate the second method, employing only the time allocation output with the identity smoothness weight. The first method fails to adapt the trajectory, leading to diverging tracking errors when waypoints are randomly shifted. In contrast, the second method can update the trajectory to remain feasible, yet it results in a longer trajectory time compared to the MFRL-trained model.
We further apply the training in a larger space ($L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$\$]) using scaled waypoint sequence data. In the wider space, trajectories could attain higher linear speeds, resulting in testing trajectories with an average speed of 5.5 $\mathrm{m}\mathrm{/}\mathrm{s}$ and a maximum of 15.1 $\mathrm{m}\mathrm{/}\mathrm{s}$, as well as an average acceleration of 0.7 g and a maximum of 2.9 g. As shown in Table 5, our trained model effectively updates trajectories while maintaining tracking accuracy, achieving even greater time reduction compared to the smaller space. In smaller spaces, tracking error bounds are typically determined by yaw tracking error due to the need for agile turns in confined areas. In contrast, in larger spaces, position tracking error becomes more prominent, as there is ample room for linear acceleration and reaching the vehicle’s maximum speed, providing our MFRL model with additional opportunities for optimization.
@@ 5.3 Analysis of MFRL Policy
Comparing the utilization factor between baseline trajectories and those from the trained model in Figure 13 reveals that the trained policy reduces the variance of the utilization factor and increases its mean. This indicates that MFRL trains the policy to maintain motor utilization at an optimal level, efficiently utilizing the vehicle’s capacity while ensuring feasibility. As shown in the right plot of Figure 13, in cases where the initial minimum snap trajectory has low motor utilization, the trained model achieves significantly greater reductions (about 12 % faster) in flight time. Furthermore, the time reduction’s variation associated with motor utilization is smaller compared to other trajectory features, highlighting their strong correlation. This reduced variation demonstrates how the trained policy consistently improves performance based on the remaining optimization margin, as quantified by motor usage. Table 6, presenting data from the default space size, and Table 7, from a larger space, compare parameters such as speed, acceleration, and motor utilization between minimum-snap and MFRL trajectories. Both tables reveal that, while average speeds and accelerations are comparable, MFRL trajectories show higher maximums and a significantly greater average motor utilization with lower variance.
Table 6: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL) ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$\$]).
@@
Table 7: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) and trajectories generated by the trained policy (MFRL) ($L_{\text{space}}=[20$\mathrm{m}$,20$\mathrm{m}$,4$\mathrm{m}$\$]).
@@
Utilizing this transformation, we analyze the performance improvements achieved through multi-fidelity evaluations and RL. By scaling the time allocation with the scale factor, we can observe trends between the scaled trajectory time and tracking error. Figure 17 compares these trends between the pretrained policy and the MFRL policy. Additionally, we train the same policy model using only low-fidelity evaluations and compare these results with those from the pretrained and RL policies developed under low-fidelity conditions. Fine-tuning the model with RL shifts this trend leftward, indicating that the planning can generate faster trajectories with similar tracking errors. Employing multi-fidelity evaluations further enhances these trends, offering improvements over training exclusively with low-fidelity models.
Figure 17: Comparison of policies trained with a multi-fidelity model: the pretrained policy (MF-MS) and the MFRL policy (MF-RL). Additionally, comparisons include the pretrained policy (LF-MS) and RL policy (LF-RL) that are trained only with the low-fidelity model. The y-axis represents the relative time reduction compared to the unscaled trajectory output from the pretrained model, which is trained using multi-fidelity evaluation.
@@ 5.4 Real-World Experiment
0.0 $\mathrm{m}$ /   0 $\mathrm{\SIUnitSymbolDegree}$ 0.14 $\pm$ 0.03 $\mathrm{m}$ 13.3 $\pm$ 1.5 $\mathrm{\SIUnitSymbolDegree}$ 0.12 $\pm$ 0.02 $\mathrm{m}$ 11.4 $\pm$ 3.8 $\mathrm{\SIUnitSymbolDegree}$ 1.69 % 0.13 $\pm$ 0.02 $\mathrm{m}$ 12.9 $\pm$ 3.7 $\mathrm{\SIUnitSymbolDegree}$ 4.26 %
1.0 $\mathrm{m}$ / 15 $\mathrm{\SIUnitSymbolDegree}$ 0.12 $\pm$ 0.02 $\mathrm{m}$ 14.8 $\pm$ 9.1 $\mathrm{\SIUnitSymbolDegree}$ 0.13 $\pm$ 0.05 $\mathrm{m}$ 14.6 $\pm$ 9.1 $\mathrm{\SIUnitSymbolDegree}$ 2.30 % 0.13 $\pm$ 0.04 $\mathrm{m}$ 16.3 $\pm$ 9.5 $\mathrm{\SIUnitSymbolDegree}$ 4.48 %
2.0 $\mathrm{m}$ / 30 $\mathrm{\SIUnitSymbolDegree}$ 0.15 $\pm$ 0.07 $\mathrm{m}$ 23.3 $\pm$ 20.2 $\mathrm{\SIUnitSymbolDegree}$ 0.12 $\pm$ 0.02 $\mathrm{m}$ 15.6 $\pm$ 13.0 $\mathrm{\SIUnitSymbolDegree}$ 2.68 % 0.13 $\pm$ 0.02 $\mathrm{m}$ 14.7 $\pm$ 7.1 $\mathrm{\SIUnitSymbolDegree}$ 3.85 %
3.0 $\mathrm{m}$ / 45 $\mathrm{\SIUnitSymbolDegree}$ 0.19 $\pm$ 0.16 $\mathrm{m}$ 29.3 $\pm$ 25.9 $\mathrm{\SIUnitSymbolDegree}$ 0.14 $\pm$ 0.05 $\mathrm{m}$ 19.4 $\pm$ 18.2 $\mathrm{\SIUnitSymbolDegree}$ 1.50 % 0.13 $\pm$ 0.04 $\mathrm{m}$ 17.7 $\pm$ 16.8 $\mathrm{\SIUnitSymbolDegree}$ 2.18 %
Table 9: Comparison of parameters including speed ($v$), acceleration ($a$), and motor utilization factor ($U_{\text{motor}}$) between minimum-snap trajectories (MinSnap) trajectories generated by the policy trained in simulation (MFRL (sim)), and trajectories generated by the policy trained with real-world experiments (MFRL (real)) ($L_{\text{space}}=[9$\mathrm{m}$,9$\mathrm{m}$,3$\mathrm{m}$\$]).
$v_{\text{avg}}$ $v_{\text{max}}$ $a_{\text{avg}}$ $a_{\text{max}}$ $U_{\text{motor}}$
We further evaluate the proposed algorithm by training the planning policy in a real-world scenario and testing the policy on a dataset of 50 randomly generated waypoint sequences, each with trajectory times $T^{\text{MS}}$ determined from real-world flight experiments. The reward estimator is trained using three different fidelity levels, including real-world flight experiments. Table 8 presents a comparison of the tracking error and the time reduction ($\Delta$ Time) achieved by policy models trained in simulated environments (using two fidelity levels) and real-world scenarios (utilizing three fidelity levels). The position and yaw tracking errors, $\bar{r}^{\text{err}}$ and $\bar{\psi}^{\text{err}}$, are calculated from the reference pose and the motion capture system measurements obtained while executing the trajectory with actual aircraft:
$\bar{r}^{\text{err}}$ $=\max_{t}\>\lVert p_{r}(t)-r_{\text{mocap}}(t)\rVert,$
$\bar{\psi}^{\text{err}}$ $=\max_{t}\>|p_{\psi}(t)-\psi_{\text{mocap}}(t)|$ (36)
where $r(t)$ and $\psi(t)$ are the reference position and yaw, respectively, and $r_{\text{mocap}}(t)$ and $\psi_{\text{mocap}}(t)$ represent the position and yaw measured using the motion capture system, as defined in (17). Notably, when waypoints deviate, both the models trained with MFRL in simulation and real-world experiments adapt the trajectory better than the baseline method. The model trained with real-world experiments further optimizes the trajectory to make it even faster than the policy only trained in simulation. Table 9 compares speed, acceleration, and motor utilization between the baseline trajectories and the trajectories from the trained policies. Similar to the simulated results, while average speeds and accelerations are comparable, MFRL trajectories show higher maximums and a significantly greater average motor utilization.
@@
Additionally, we assess the trained model’s performance in an online planning scenario, where dynamically shifting waypoints necessitate real-time trajectory adaptation within the environment shown in Figure 18. We put the waypoints in the middle of two gates and continuously moving the gates to update the position of the waypoints. We commit a portion of the first trajectory segment, taking into account inference and communication delays, while adjusting the time allocation for the remainder. The time allocation between waypoints is determined by the trained policy, and we obtain the actual polynomial trajectory through quadratic programming. The trained policy is executed on the Titan Xp GPU in the host computer and communicates with the vehicle’s microcontroller to update the trajectory. By compiling the model with TensorRT, we significantly enhance the inference speed. This enables our trained model to adjust the trajectory in under 2 milliseconds, allowing real-time trajectory adaptation—a marked contrast to the baseline minimum-snap method, which takes several minutes for trajectory generation. Since the planning policy comprises simple MLPs, the inference time remains consistent even on embedded GPUs, averaging 3 milliseconds when executed on Jetson Orin. Table 10 presents a comparison of inference times between the baseline method and our trained model. The baseline method’s line search procedure significantly increases computation time, especially when using real-world flight experiment data where evaluating multiple speed profiles for a trajectory can take several minutes. Generally, the minimum snap method is used with the lowest fidelity line search, but even this approach requires around 400 ms, substantially longer than our method. Our proposed method streamlines this by substituting the line search with policy inference, making it suitable for real-time applications. All computation times are measured on an Intel Xeon CPU with 28 cores in the host computer, except for the policy model, which is evaluated on the aforementioned Titan Xp GPU in the same machine. Figure 19 compares the inference times based on the number of waypoints. Even with a small number of waypoints, the policy inference time is substantially lower, requiring far less computation time. The supplementary video includes demonstrations of the real-world experimental results.